The rectangle is 33 feet long and 25 feet wide if the gardener wants to build a fence around the garden how many feet of fence are required
Answer by ikleyn(52780)

.
The rectangle is 33 feet long and 25 feet wide. If the gardener wants to build a fence around the garden, how many feet of fence are required ?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Two times 33 feet plus two times 25 feet, in total

2*(33 + 25) = 2*58 = 116 feet.